Pore O2 diffuses thro an apparatus in 224 sec., whereas mixture of O2 and other gas containing 80 % O2 diffuses from same in 234 s. What is Molecukar weight of gas?

Grade:11

1 Answers

askiitian.expert- chandra sekhar
14 years ago

Hi Vishrnat,

r1/r2=234/224=√M2/M1

M1=32

M2=0.8*32+0.2*M

M2/M1=1.0906

0.8 + 0.2*M/32=1.0906

M= 0.29*32/0.2 ≈ 47